Given W is a uniformly distributed random variable with mean 33 and variance 3. Using R programming, apply the suitable coding to identify answer in (a) and (b) 


(a) probability density function for W 

(b) cumulative distribution function for W 


1
Expert's answer
2022-02-16T07:09:03-0500

(a) probability density function for W 

we define the function dnorm: which evaluates the normal probability density with a given mean or

standard deviation

The coding is as below

dnorm( x, mean = 33, sd = 31/2, log = FALSE )


(b) cumulative distribution function for W 

we define the function pnorm: which evaluates the cumulative distribution function for a normal

distribution.

The coding is as below

pnorm( q, mean = 33, sd = 31/2, lower.tail =TRUE, log.p = FALSE )